For much of the 20th century, new car prices rose at an average annual rate of 5.89%. Given a beginning new car price of $900, how long did it take for the average new car price to rise to $27,865? Please round to the nearest year.
Answer #1

The initial price of the car price is P0 = $900.

The final price of the car price is Pn = $27,865.

Where, n is the number of years in which this increase in the average car prices happened.

Given annual rate, r at which the car prices rose is 5.89%.

We know that, Final value = Initial value * ( 1 + r% )n

Therefore, on substituting in the above formula we get:

27,865 = 900 * (1+0.0589)n

27,865 / 900 = (1.0589)n

30.96 = (1.0589)n

Applying log on both sides.

log(30.96) = log(1.0589n)

We know that logab = b * loga

Therefore, log(30.96) = n * log(1.0589)

Here, n = log(30.96) / log(1.0589)

Substituting the logarithmic values in the above equation:

n = 1.491 / 0.0248

n = 60

Therefore, it took 60 years in 20th century for the average car prices to rise from $900 to $27,865.